OpenStudy (thatgirl21):

Please help! Given the chart below: x -6 -1 3 5 8 13 g(x) 22 7 -5 -11 -20 -35 answer each of the following a.) find a formula for this set of data b.) what is the value of g(-21)? c.) find the value of k such that g(k)=k-12

OpenStudy (thatgirl21):

for a I don't know if I'm doing it right. I found the slope to be -3 because m=(22-7)/(-6+1) m=15/-5 m=-3

OpenStudy (thatgirl21):

now I think I do y-y1=m(x-x1)? not really sure